Find the general solution.

2 16. 7(t) = -3 1 - 2 0 - 1 -1 0 (t) 0
Answer #1

16. ne (t) -3 2. 이 다 - O al4-0 2 o let net kyle) 72(1) Azet) CU alt) (t) ܪܳܟ ng (6) from ny na ta :) 21 maz 57 es 324 t. 207 (-3-20 (a+a?) +2 2{1+2(1-2)}=0 -32-22-322-23 Ô -42= 0 -23 - 482-71-6--0 +422 +72+ 6 = 0 (A +2) C a2 +22+3)=0 93 71 2= -2, 2From lot equation of 6 we have, 2x3 3x + 344 = -2922tét (Bcostit +C stoort) tet (i sindetto CoS+24) + 386 24 +38 (8695124 +Cgeneral so the gener al solution is 24(H) = Het tet | BCOS124 + e sinda 1024) cot 28(t) = -AQ2E+ Bet sincet COS<2l+pet A3 (t)